A heavy lineman tackles a light running back. Who experiences the greater acceleration?

the lineman
they both experience the same acceleration
the running back

Flag this Question

Question 71 pts

If you are 5.5 slugs on earth, how many slugs would you be on the moon?

more
less
the same

Flag this Question

Question 81 pts

If a book weighs 10 lbs, what is the total force on it while it is in Free Fall?

10 lbs
0 lbs
20 lbs
32 lbs

Flag this Question

Question 91 pts

You fire a gun at 600 m/s in a horizontal direction. One second later, what are the horizontal and vertical components of the speed of this gun, respectively? (Ignore air friction)

500, 9.8
600, 9.8
600, 0
0, 0
Answer #1

1. As the lineman is running with the bag, so bag is also moving with man, so bag is also moving with the same acceleration as that of man. If their acceleration are different, they can not move to gether.

2.Slug is only the mass , so it does not change with change in planet, rather its weight that is pound will change if we move from earth to moon, on moon, its weight will be around 1/6 of earth as gravity of moon is 1/6 of earth.

3.DUring free fall, the force will be only gravitational force, so it will be 10 pound.

4. In horizontal direction, the velocity does not change as there is no force in horizontal direction, so its horizontal velocity will be 600 m/s.

In vertical direction, gravity acts downwarfds, so its velocity in vertically downward direction will increase gradually.

Using v = u + gt, where u = 0, g = 9.8

v = 0+ 9.8 = 9.8.